# Accelerating the Convergence of Coupled Cluster Calculations of the Homogeneous Electron Gas Using Bayesian Ridge Regression

### v2

@@ Abstract
The homogeneous electron gas is a system which has many applications in chemistry and physics. However, its infinite nature makes studies at the many-body level complicated due to long computational run times. Because it is size extensive, coupled cluster theory is capable of studying the homogeneous electron gas, but it still poses a large computational challenge as the time needed for precise calculations increases in a polynomial manner with the number of particles and single-particle states. Consequently, achieving convergence in energy calculations becomes challenging, if not prohibited, due to long computational run times and high computational resource requirements. This paper develops the sequential regression extrapolation (SRE) to predict the coupled cluster energies of the homogeneous electron gas in the complete basis limit using Bayesian ridge regression to make predictions from calculations at truncated basis sizes. Using the SRE method we were able to predict the coupled cluster doubles energies for the electron gas across a variety of values of N and $r_{s}$, for a total of 70 predictions, with an average error of 4.28x10-4 Hartrees while saving 88.9 hours of computational time. The SRE method can accurately extrapolate electron gas energies to the complete basis limit, saving both computational time and resources. Additionally, the SRE is a general method that can be applied to a variety of systems, many-body methods, and extrapolations.
@@ Introduction
Singal and Das Singal and Das (1973) were the first to study the electron gas using a CC approach. The approximation they used is similar to the Brueckner Hartree-Fock method Haftel and Tabakin (1970), and did not properly include particle-hole ring diagrams. Later, Freeman performed CC calculations Freeman (1977) in which only ring diagrams and their exchange parts were retained in a CC doubles approximation. The results of both Singal and Das Singal and Das (1973) and Freeman Freeman (1977) compared well with dielectric-function approaches. These calculations of the three-dimensional electron gas were improved by Bishop and Lührmann Bishop and L”uhrmann (1978, 1982). Bishop and Lührmann derived a CC SUB2 approximation (also called CC doubles (CCD) Shavitt and Bartlett (2009b)) for the electron gas, extended with some ladder contributions from higher-order amplitudes. As the authors described in Ref. Bishop and L”uhrmann (1978), the CCD approximation contains many more diagrammatic classes than partial-summation techniques derived from perturbation theory. The CCD approximation takes account of particle-particle and hole-hole ladders, particle-hole ring diagrams including exchange terms, and many other diagrams to infinite order in perturbation theory. Recently, Shepherd et al. Shepherd et al. (2012b); Shepherd and Grüneis (2013); Shepherd et al. (2014a, b) have taken up again the CC effort for the three-dimensional electron gas. The CC calculations for the electron gas from the 1970s and 1980s were all done at the thermodynamic limit Singal and Das (1973); Freeman (1977); Bishop and L”uhrmann (1978, 1982). Unfortunately, singularities related to the Coulomb two-body interaction may complicate numerical calculations at the thermodynamic limit. Shepherd et al. approximate the electron gas using finite cubic boxes Shepherd et al. (2012b); Shepherd and Grüneis (2013); Shepherd et al. (2014a, b). However, to make CC calculations feasible in the HEG using finite lattices, truncations are necessary, thereby introducing errors that diminish accuracy. The truncation in the basis of single-particle states ($M$) leads to basis incompleteness errors while truncating the number of electrons ($N$) and the HEG volume ($\Omega$) introduces finite-size errors. One way to reduce these errors is by performing CC calculations at high values of $M$ and $N$, such that the calculations have converged. However, this approach demands extensive computational resources and time. Since basis incompleteness errors and finite size errors arise from different truncations, they can be addressed independently Shepherd et al. (2012a); Shepherd (2016). This work focuses on removing the basis truncation errors from CC doubles calculations of the HEG using machine learning.
@@
Coupled cluster theory starts with the exponential ansatz that states a many-body wavefunction can be created by applying an exponential cluster operator to the Fermi vacuum state, shown in Eq. (1). The cluster operator consists of N excitation operators, where N is the number of electrons or particles in the system. However, for practical calculations, the number of terms in the cluster operator is often truncated to a set level Taube and Bartlett (2008); Kucharski and Bartlett (1998); Bartlett et al. (1990); Raghavachari et al. (1989); He et al. (2001); Urban et al. (1985). In this work, the focus is on the doubles level (CCD), which approximates the cluster operator with just the two-particle two-hole excitation operator (Eq. (2)).
@@
$\psi_{j}(\vec{r},\sigma)=\sqrt{\frac{1}{\Omega}}\ e^{i\vec{k}j\cdot\vec{r}}\ \delta{\sigma_{j},\sigma},$ (6)
@@
(a) $\Delta E_{CC}$ per electron plotted against $N$ with calculations performed at four different values of $M$. As $M$ increases, the error in the calculation decreases, but the run time drastically increases.
(b) The run times (measure in node hours) for CCD calculations of the HEG as a function of M and at three different values of N.
@@
(d) A plot of $\Delta E_{CC}$ and $\Delta E_{MBPT}$ versus $M$ for a HEG system with $N=342$ and rs=0.5. The converged correlation energies calculated at $M=6,142$ are shown with dashed lines for each method.
[…]
(f) A plot of the ratio of $\Delta E_{CC}$ to $\Delta E_{MBPT}$ as a function of $M$ for $N=114$, $342$, and $514$ and rs=0.5. As $M$ increases, the ratio (or the slope of the line in Figure 1(c) becomes increasingly constant. However, it is important to note that the run time increases drastically as $M$ increases.
Figure 1: An analysis of the relationship between $\Delta E_{CC}$ and $\Delta E_{MBPT}$ for calculations on the HEG at various $M$, $N$, and rs.
In this work, the CCD and MBPT2 correlation energies will be denoted as $\Delta$ECC,M and $\Delta$EMBPT,M, respectively, to show the number of single-particle states in the calculations. Figure 1(a) shows a plot of $\Delta E_{CC}$ per particle for all magic numbers between two and 514 (inclusive), with the CC calculations performed at four different values of $M$ and at $r_{s}$ = 0.5. The calculations performed at $M=514$ (the red line in the plot) are the least accurate since the data was generated with the lowest value of $M$. When $M=6,142$, the correlation energies are converged (to $10^{-7}$ Hartrees), making this data set (plotted in purple) the most accurate. However, as $M$ increased, the run time and computational resources of the calculation drastically increased as well.
@@
Therefore, this paper aims to predict the converged $\Delta E_{CC}$ for the HEG using only data calculated at smaller values of $M$, where the computational time and resource requirements are less significant. This paper will use $\Delta E_{CC,6142}$ as the actual, or most accurate result, and $\Delta E_{CC}$ has been confirmed to have converged at this point, meaning that the correlation energies have converged to within $10^{-7}$ Hartrees when adding additional single-particle states to the calculations.
Figure 1(d) shows the convergence of $\Delta E_{CC}$ as a function of $M$ for a HEG with $N=342$ and rs=0.5. Though the apparent route would be to attempt to predict the converged correlation energy (plotted with a red dashed line) using the converging correlation energies from smaller $M$ calculations, in practice, predictions using this method do not have an acceptable accuracy. However, if we plot $\Delta E_{MBPT}$ for the same system in addition to $\Delta E_{CC}$, we see that they follow a very similar pattern. Since $\Delta E_{MBPT}$ can be generated quickly, the goal of this work is to create a model through which the converged $\Delta E_{CC}$ can be predicted from $\Delta E_{MBPT}$. Figure 1(e) is a transformation of Figure 1(d), where $\Delta E_{CC}$ are plotted as a function of $\Delta E_{MBPT}$. Each point represents a calculation performed with the same value of $M$. When the data is formatted in this way, $M$ increases from right to left. As the graph moves from right to left, it becomes increasingly linear (i.e., the ratio $\Delta E_{CC}/\Delta E_{MBPT}$ becomes constant as $M$ increases).
We can see this trend more clearly if we plot $\Delta E_{CC}/\Delta E_{MBPT}$ for various $M$ values and $N=114$, $342$, and $514$ with rs = 0.5 in Figure 1(f). For each of these data sets at larger values of $M$, the ratio leveled off to a constant, or near constant, value. Therefore the goal of the research is to be able to predict the constant value of $\Delta E_{CC}/\Delta E_{MBPT}$, which occurs at larger values of $M$ by extrapolating from data points collected at lower values of $M$. We will use this pattern to create the input to a method created by the authors known as sequential regression extrapolation (SRE), a regression-based machine learning algorithm used to accurately extrapolate small data sets generated from many-body codes.
Both $\Delta E_{CC}$ and $\Delta E_{MBPT}$ converge as M increases. Therefore, there must be a large value of M where their ratio becomes a constant:
$\frac{\Delta E_{CC,Large\ M}}{\Delta E_{MBPT,Large\ M}}\longrightarrow constant.$ (7)
@@
$y=\frac{\Delta E_{CC,M}}{\Delta E_{MBPT,M}}=\frac{\Delta E_{CC,M_{1}}}{\Delta E_{MBPT,M_{1}}},\frac{\Delta E_{CC,M_{2}}}{\Delta E_{MBPT,M_{2}}},\frac{\Delta E_{CC,M_{3}}}{\Delta E_{MBPT,M_{3}}},...,\frac{\Delta E_{CC,M_{15}}}{\Delta E_{MBPT,M_{15}}}$ (8)
Next, a ridge regression algorithm with Bayesian hyperparameter tuning is trained on this data set Murphy (2013); MacKay (1992); Tipping (2001). However, the way this algorithm is trained is unique because there is no x component of the data set. Instead, the input to the algorithm is two sequential ratios, taking at neighboring values of M, and the output is the ratio at the next value of M. This method of data formatting combined with a regression algorithm gives rise to the SRE algorithm.
$f_{Ridge}(\frac{\Delta E_{CC,k-2}}{\Delta E_{MBPT,k-2}},\frac{\Delta E_{CC,k-1}}{\Delta E_{MBPT,k-1}})=\frac{\Delta E_{CC,k}}{\Delta E_{MBPT,k}}$ (9)
The SRE algorithm is then used to extrapolate this data set to many points until the ratio of correlation energies has converged (for 100 points in total). This value can then be taken as the slope of the graph created with $\Delta E_{CC}$ plotted as a function of $\Delta E_{MBPT}$ at high M (Figure 1(e)).
$\lim_{k\to\infty}\frac{\Delta E_{CC,k}}{\Delta E_{MBPT,k}}=slope=m$ (10)
Finally, $\Delta E_{MBPT,6142}$ is generated, a process that takes less than one second (not including the time to generate the matrix elements). $\Delta E_{MBPT,6142}$ is multiplied by the slope, m, to approximate $\Delta E_{CC,6142}$.
$m\Delta E_{MBPT,6142}\approx\Delta E_{CC,6142}$ (11)
@@
The time needed to train the SRE algorithm and predict the converged correlation energy is on the order of milliseconds, so the primary time savings will come from calculating the training data versus the time needed to perform the calculation at $M=6142$. When the number of electrons $N$ is small, the time savings is also small, only 0.024 node hours for $N=66$ and $r_{s}$ = 0.5. However, the time savings grows quickly as $N$ increases. For example, at $N=246$ and $r_{s}$ = 0.5, using the SRE method saves 2.05 node hours, and by N=514 and $r_{s}$ = 0.5, the SRE time savings is 5.88 node hours. Therefore, the SRE method can drastically reduce the run times of predicting the converged correlation energies for the HEG system, especially for studies requiring many values of rs and $N$. Additionally, the total time saved from just generating the training data compared to performing all calculations at $M=6142$ for all data points shown in Figure 2 is 88.9 node hours.
The amount of RAM needed to generate the training points with the highest $M$ value ($M=2090$) is only 33.15 GB at 514 electrons (represented the largest combination of $N$ and $M$ across all of the training data sets) while it takes 106.48 GB of RAM just to generate the fully converged coupled cluster correlation energy at 66 electrons (one of the smaller numbers of electrons studied). At its worst, calculating the fully converged CCD correlation energy at 514 electrons (the highest number of electrons in this study) took 444.48 GB of RAM. Thus, using the SRE method and only needing to generate a small amount of training data at low values of M saves significant computational time and resources when compared to performing the fully converged calculations at a high value of $M$.
While a reduced run time and lessened computational resources are important, an even more important factor is the accuracy of the predictions. In Figure 2, the predicted converged $\Delta E_{CC}$ per electron is plotted with $\Delta E_{CC,6142}/N$ for a variety of values of $N$ and rs. We limit ourselves to values of $r_{s}<1.0$ as we are interested in studying a high-density electron gas Carr and Maradudin (1964); Sawada (1957). The solid lines graph the fully converged calculations performed at $M=6142$, the triangular markers plot the SRE prediction for the converged CCD correlation energy per electron at each combination of $N$ and $r_{s}$, and the shaded regions show the uncertainty of the ML prediction, computed by the Bayesian ridge regression algorithm. Each color in the plot represents a different value of $r_{s}$, as shown in the legend of Figure 2.
Figure 2: The results from performing an SRE extrapolation for various numbers of electrons and values of rs. Results are plotted against $\Delta E_{CC,6142}$ and the uncertainties on the predictions are also shown with shaded regions.
@@
Table 1: A quantitative summary of the SRE prediction accuracy. The first column is the Wigner-Seitz radius, the second the average error (in Hartrees) for each value of $r_{s}$, and the third column the average uncertainty on the predictions, taken from the Bayesian ridge regression algorithm.

### v3

@@ Abstract
The homogeneous electron gas is a system that has many applications in chemistry and physics. However, its infinite nature makes studies at the many-body level complicated due to long computational run times. Because it is size extensive, coupled cluster theory is capable of studying the homogeneous electron gas, but it still poses a large computational challenge as the time needed for precise calculations increases in a polynomial manner with the number of particles and single-particle states. Consequently, achieving convergence in energy calculations becomes challenging, if not prohibited, due to long computational run times and high computational resource requirements. This paper develops the sequential regression extrapolation (SRE) to predict the coupled cluster energies of the homogeneous electron gas in the complete basis limit using Bayesian ridge regression and many-body perturbation theory correlation energies to the second order to make predictions from calculations at truncated basis sizes. Using the SRE method we were able to predict the coupled cluster doubles energies for the electron gas across a variety of values of N and $r_{s}$, for a total of 70 predictions, with an average error of 5.20x10-4 Hartrees while saving 88.9 hours of computational time. The SRE method can accurately extrapolate electron gas energies to the complete basis limit, saving both computational time and resources. Additionally, the SRE is a general method that can be applied to a variety of systems, many-body methods, and extrapolations.
@@ Introduction
Singal and Das Singal and Das (1973) were the first to study the electron gas using a CC approach. The approximation they used is similar to the Brueckner Hartree-Fock method Haftel and Tabakin (1970), and did not properly include particle-hole ring diagrams. Later, Freeman performed CC calculations Freeman (1977) in which only ring diagrams and their exchange parts were retained in a CC doubles approximation. The results of both Singal and Das Singal and Das (1973) and Freeman Freeman (1977) compared well with dielectric-function approaches. These calculations of the three-dimensional electron gas were improved by Bishop and Lührmann Bishop and Lührmann (1981); Bishop and L”uhrmann (1982). Bishop and Lührmann derived a CC SUB2 approximation (also called CC doubles (CCD) Shavitt and Bartlett (2009a)) for the electron gas, extended with some ladder contributions from higher-order amplitudes. As the authors described in Ref. Bishop and Lührmann (1981), the CCD approximation contains many more diagrammatic classes than partial-summation techniques derived from perturbation theory. The CCD approximation takes account of particle-particle and hole-hole ladders, particle-hole ring diagrams including exchange terms, and many other diagrams to infinite order in perturbation theory. Recently, Shepherd et al. Shepherd et al. (2012a); Shepherd and Grüneis (2013); Shepherd et al. (2014a); Bishop and Lührmann (1982) have taken up again the CC effort for the three-dimensional electron gas. The CC calculations for the electron gas from the 1970s and 1980s were all done at the thermodynamic limit Singal and Das (1973); Freeman (1977); Bishop and Lührmann (1981); Bishop and L”uhrmann (1982). Unfortunately, singularities related to the Coulomb two-body interaction may complicate numerical calculations at the thermodynamic limit. Shepherd et al. approximate the electron gas using finite cubic boxes Shepherd et al. (2012a); Shepherd and Grüneis (2013); Shepherd et al. (2014a); Bishop and Lührmann (1982). However, to make CC calculations feasible in the HEG using finite lattices, truncations are necessary, thereby introducing errors that diminish accuracy. The truncation in the basis of single-particle states ($M$) leads to basis incompleteness errors while truncating the number of electrons ($N$) and the HEG volume ($\Omega$) introduces finite-size errors. One way to reduce these errors is by performing CC calculations at high values of $M$ and $N$, such that the calculations have converged. However, this approach demands extensive computational resources and time. Since basis incompleteness errors and finite size errors arise from different truncations, they can be addressed independently Shepherd et al. (2012a); Shepherd (2016). This work focuses on removing the basis truncation errors from CC doubles calculations of the HEG using machine learning.
@@
Coupled cluster theory starts with the exponential ansatz that states a many-body wavefunction can be created by applying an exponential cluster operator to the Fermi vacuum state, shown in Eq. (1). The cluster operator consists of N excitation operators, where N is the number of electrons or particles in the system. However, for practical calculations, the number of terms in the cluster operator is often truncated to a set level Taube and Bartlett (2008); Kucharski and Bartlett (1998); Bartlett et al. (1990); Raghavachari et al. (1989); He et al. (2001); Urban et al. (1985). In this work, the focus is on the doubles level (CCD), which approximates the cluster operator with just the two-particle two-hole excitation operator (Eq. (2)). Since the homogeneous electron gas is an infinite matter system, the terms which result from coupled cluster singles are zero due to momentum conservation. Thus performing CCD calculations here are equivalent to performing coupled cluster singles and doubles (CCSD) calculations.
@@
In the above equation, $E_{REF}$ is the reference energy obtained from Hartree-Fock theory, $\epsilon_{p}$ is the energy of state p, and $\langle ij|\hat{v}_{2}|ab\rangle$ is the two-body operator between the filled states i and j, and the empty states a and b.
@@
$\psi_{j}(\vec{r},\sigma)=\sqrt{\frac{1}{\Omega}}\ e^{i\vec{k}j\cdot\vec{r}}\ \delta_{\sigma_{j},\sigma},$ (6)
@@
(a) $\Delta E_{CCD}$ per electron plotted against $N$ with calculations performed at four different values of $M$. As $M$ increases, the error in the calculation decreases, but the run time drastically increases.
(b) The run times (measured in node hours) for CCD calculations of the HEG as a function of M and at three different values of N.
@@
(d) A plot of $\Delta E_{CCD}$ and $\Delta E_{MBPT2}$ versus $M$ for a HEG system with $N=342$ and rs=0.5. The converged correlation energies calculated at $M=6,142$ are shown with dashed lines for each method.
(e) Figure 1(d) is transformed such that $\Delta E_{CCD}$ are plotted as a function of $\Delta E_{MBPT2}$ (such that each point has the same $M$). The resulting plot has a linear relationship that becomes stronger to the left (as $M$ increases).
(f) A plot of the ratio of $\Delta E_{CCD}$ to $\Delta E_{MBPT2}$ as a function of $M$ for $N=114$, $342$, and $502$ and rs=0.5. As $M$ increases, the ratio (or the slope of the line in Figure 1(c) becomes increasingly constant. However, it is important to note that the run time increases drastically as $M$ increases.
Figure 1: An analysis of the relationship between $\Delta E_{CCD}$ and $\Delta E_{MBPT2}$ for calculations on the HEG at various $M$, $N$, and rs.
In this work, the CCD and MBPT2 correlation energies will be denoted as $\Delta$ECCD,M and $\Delta$EMBPT2,M, respectively, to show the number of single-particle states in the calculations. Figure 1(a) shows a plot of $\Delta E_{CCD}$ per particle for all magic numbers between two and 502(inclusive), with the CC calculations performed at four different values of $M$ and at $r_{s}$ = 0.5. The calculations performed at $M=514$ (the red line in the plot) are the least accurate since the data was generated with the lowest value of $M$. When $M=6,142$, the correlation energies are converged (to $10^{-7}$ Hartrees), making this data set (plotted in purple) the most accurate. However, as $M$ increased, the run time and computational resources of the calculation drastically increased as well.
@@
Therefore, this paper aims to predict the converged $\Delta E_{CCD}$ for the HEG using only data calculated at smaller values of $M$, where the computational time and resource requirements are less significant. This paper will use $\Delta E_{CCD,6142}$ as the actual, or most accurate result, and $\Delta E_{CDC}$ has been confirmed to have converged at this point, meaning that the correlation energies have converged to within $10^{-7}$ Hartrees when adding additional single-particle states to the calculations.
Figure 1(d) shows the convergence of $\Delta E_{CCD}$ as a function of $M$ for a HEG with $N=342$ and rs=0.5. Though the apparent route would be to attempt to predict the converged correlation energy (plotted with a red dashed line) using the converging correlation energies from smaller $M$ calculations, in practice, predictions using this method do not have an acceptable accuracy. However, if we plot $\Delta E_{MBPT2}$ for the same system in addition to $\Delta E_{CCD}$, we see that they follow a very similar pattern. Since $\Delta E_{MBPT2}$ can be generated quickly, the goal of this work is to create a model through which the converged $\Delta E_{CCD}$ can be predicted from $\Delta E_{MBPT2}$. Figure 1(e) is a transformation of Figure 1(d), where $\Delta E_{CCD}$ are plotted as a function of $\Delta E_{MBPT2}$. Each point represents a calculation performed with the same value of $M$. When the data is formatted in this way, $M$ increases from right to left. As the graph moves from right to left, it becomes increasingly linear (i.e., the ratio $\Delta E_{CCD}/\Delta E_{MBPT2}$ becomes constant as $M$ increases).
We can see this trend more clearly if we plot $\Delta E_{CCD}/\Delta E_{MBPT2}$ for various $M$ values and $N=114$, $342$, and $502$ with rs = 0.5 in Figure 1(f). For each of these data sets at larger values of $M$, the ratio leveled off to a constant, or near constant, value. Therefore the goal of the research is to be able to predict the constant value of $\Delta E_{CCD}/\Delta E_{MBPT2}$, which occurs at larger values of $M$ by extrapolating from data points collected at lower values of $M$. We will use this pattern to create the input to a method created by the authors known as sequential regression extrapolation (SRE), a regression-based machine learning algorithm used to accurately extrapolate small data sets generated from many-body codes.
Both $\Delta E_{CCD}$ and $\Delta E_{MBPT2}$ converge as M increases. Therefore, there must be a large value of M where their ratio becomes a constant:
$\frac{\Delta E_{CCD,Large\ M}}{\Delta E_{MBPT2,Large\ M}}\longrightarrow constant.$ (7)
@@
$y=\frac{\Delta E_{CCD,M}}{\Delta E_{MBPT2,M}}=\frac{\Delta E_{CCD,M_{1}}}{\Delta E_{MBPT2,M_{1}}},\frac{\Delta E_{CCD,M_{2}}}{\Delta E_{MBPT2,M_{2}}},\frac{\Delta E_{CCD,M_{3}}}{\Delta E_{MBPT2,M_{3}}},...,\frac{\Delta E_{CCD,M_{15}}}{\Delta E_{MBPT2,M_{15}}}$ (8)
Next, a ridge regression algorithm with Bayesian hyperparameter tuning is trained on this data set Murphy (2013); MacKay (1992); Tipping (2001). The form of the Bayesian ridge regression, $f_{Ridge}$ in Eq. 9, are defined in Refs. MacKay (1992) and Tipping (2001). However, the way this algorithm is trained is unique because there is no x component of the data set. Instead, the input to the algorithm is two sequential ratios, taking at neighboring values of M, and the output is the ratio at the next value of M. This method of data formatting combined with a regression algorithm gives rise to the SRE algorithm. There are, however, a few important points to note in regards to how the SRE algorithm is trained and used. First, the algorithm is only trained for a given system with set parameters; thus each data point in 2 represents a separate training and prediction of the SRE algorithm. The values of $M$ chosen for the training data are taken to correspond to closed energy shells, with the exact values of $M$ ranging between one and 15 open shells with the exact values of $M$ depending on the number of electrons (over the full energy shells which hold the electrons in the ground state). Thus, the values of $M$ used in the calculations will vary between calculations with different numbers of electrons, though changing the density does of affect the values of $M$ used in the calculation.
$f_{Ridge}(\frac{\Delta E_{CCD,k-2}}{\Delta E_{MBPT2,k-2}},\frac{\Delta E_{CCD,k-1}}{\Delta E_{MBPT2,k-1}})=\frac{\Delta E_{CCD,k}}{\Delta E_{MBPT2,k}}$ (9)
When Eq. 9 is trained using 15 training points at different values of M, the time-series like formatting results in eleven training points: (($M_{1}$, $M_{2}$), $M_{3}$), (($M_{2}$, $M_{3}$), $M_{4}$), (($M_{3}$, $M_{4}$), $M_{5}$), … . Once trained, the SRE algorithm is then used to extrapolate this data set to many points until the ratio of correlation energies has converged (for 50 points in total). This value can then be taken as the slope of the graph created with $\Delta E_{CCD}$ plotted as a function of $\Delta E_{MBPT2}$ at high M (Figure 1(e)).
$\lim_{k\to\infty}\frac{\Delta E_{CCD,k}}{\Delta E_{MBPT2,k}}=slope=m$ (10)
Finally, $\Delta E_{MBPT2,6142}$ is generated, a process that takes less than one second (not including the time to generate the matrix elements). $\Delta E_{MBPT2,6142}$ is multiplied by the slope, m, to approximate $\Delta E_{CCD,6142}$.
$m\Delta E_{MBPT2,6142}\approx\Delta E_{CCD,6142}$ (11)
@@
In the analysis of the results and their stability and also range of applicability, one should keep in mind that many-body perturbation theory and coupled cluster theories are theories which explicitly include energy denominators defined in terms of the differences between states below and above the Fermi level. When the system (and this applies to both finite and infinite systems like the electron gas studied here) becomes more dilute, which corresponds to larger values of $r_{s}$, these energy gaps become smaller and smaller and result eventually in a worsened convergence and/or unstable results as function of the number of single-particle states. For larger values of $r_{s}$ one needs more complicated correlations than those included here in order to obtain stable results. For values of $r_{s}$ larger than those used in this work, many-body perturbation theory to second order results normally in less table results as function of intermediate single-particle states, and limits thus the applicability (with its simplicity defined by second-order MBPT) of the present method.
@@
The time needed to train the SRE algorithm and predict the converged correlation energy is on the order of milliseconds, so the primary time savings will come from calculating the training data versus the time needed to perform the calculation at $M=6,142$. When the number of electrons $N$ is small, the time savings is also small, only 0.024 node hours for $N=66$ and $r_{s}$ = 0.5. However, the time savings grows quickly as $N$ increases. For example, at $N=246$ and $r_{s}$ = 0.5, using the SRE method saves 2.05 node hours, and by N=502 and $r_{s}$ = 0.5, the SRE time savings is 5.88 node hours. Therefore, the SRE method can drastically reduce the run times of predicting the converged correlation energies for the HEG system, especially for studies requiring many values of rs and $N$. Additionally, the total time saved from just generating the training data compared to performing all calculations at $M=6,142$ for all data points shown in Figure 2 is 88.9 node hours.
The amount of RAM needed to generate the training points with the highest $M$ value ($M=2,090$) is only 33.15 GB at 502 electrons (represented the largest combination of $N$ and $M$ across all of the training data sets) while it takes 106.48 GB of RAM just to generate the fully converged coupled cluster correlation energy at 66 electrons (one of the smaller numbers of electrons studied). At its worst, calculating the fully converged CCD correlation energy at 502 electrons (the highest number of electrons in this study) took 444.48 GB of RAM. Thus, using the SRE method and only needing to generate a small amount of training data at low values of M saves significant computational time and resources when compared to performing the fully converged calculations at a high value of $M$.
While a reduced run time and lessened computational resources are important, an even more important factor is the accuracy of the predictions. In Figure 2, the predicted converged $\Delta E_{CCD}$ per electron is plotted with $\Delta E_{CCD,6142}/N$ for a variety of values of $N$ and rs. We limit ourselves to values of $r_{s}<1.0$ as we are interested in studying a high-density electron gas Carr and Maradudin (1964); Sawada (1957). The solid lines graph the fully converged calculations performed at $M=6,142$, the triangular markers plot the SRE prediction for the converged CCD correlation energy per electron at each combination of $N$ and $r_{s}$, and the shaded regions show the uncertainty of the ML prediction, computed by the Bayesian ridge regression algorithm. Each color in the plot represents a different value of $r_{s}$, as shown in the legend of Figure 2.
Figure 2: The results from performing an SRE extrapolation for various numbers of electrons and values of rs. Results are plotted against $\Delta E_{CCD,6142}$ and the uncertainties on the predictions are also shown with shaded regions.
@@
Table 1: A quantitative summary of the SRE prediction accuracy. The first column is the Wigner-Seitz radius, the second the average error (in Hartrees) for each value of $r_{s}$, and the third column is the average uncertainty on the predictions, taken from the Bayesian ridge regression algorithm.